# Complex networks approach to curriculum analysis and subject integration: a case study on Physics and Mathematics

### v2

@@ I Introduction
The application of network science to education is not a novel concept and has been explored in various studies Koponen and Mäntylä (2020). These methods not only facilitate the visualization of complex relationships but also enable the quantification of indirect effects that traditional statistical methods may fail to capture. There are several studies on the coherence of content blocks through the analysis of community structures within students’ concept networks and the centrality of their nodes. Various examples of the use of concept networks are the following: Some studies about general Sayama (2017) or specific topics such as energy and electrostatics Podschuweit and Bernholt (2020); Thurn et al. (2020) use these networks to evaluate and understand students’ ideas and misconceptions. They are also used to evaluate learning processes Siew (2020) or inferential reasoning chains of physics concepts Speirs et al. (2024). In Kubsch et al. (2020), they establish a correlation between the coherence of concept networks and the practical application of knowledge, while in Schwab et al. (2017) teaching sequences are personalized by considering the complex relationships between concepts. Related to concept networks, there are also studies analyzing students’ answers to multiple choice surveys about physics concepts that provide interesting information about students’ misconceptions related to other variables Stewart et al. (2021); Wells et al. (2020); Wheatley et al. (2022). Apart from concept networks, network theory in education has been used to evaluate, for example, classmates networks and students’ academic gain Zwolak et al. (2018); Traxler et al. (2018), considering different contexts such as teaching mode (online vs. in-person) Pulgar et al. (2023) or teaching approach (active learning) Traxler et al. (2020). Similarly, authorship networks in publications within the field have also been analyzed Anderson et al. (2017), and social practices between instructors and students have been described using network theory Olsen et al. (2023).
@@ I.1 Basic concepts about complex networks
Centrality measures describe the importance of nodes in a network in different ways. There are many centralities, but in this work we are going to concentrate on three: degree, closeness and betweenness centrality. The degree is the number of links connected to a node and is the most basic centrality measure. It is a local measure that does not require knowledge of the global structure of the network. Typically, a highly connected node is considered very important or central to the network. Nodes with disproportionately large degrees in the network are referred to as hubs. In content networks, nodes with higher degrees represent content that can be related to many other contents, either due to their cross-disciplinary nature, their broad applicability, or because they combine results from different contents. Closeness is the inverse of the average length of the shortest paths that connect a node to the rest of the nodes in the network. In this way, a node that is close to the rest of the nodes will be considered central. This measure requires global knowledge of the network. In content networks, this centrality measure will be able to identify the content that can be more easily related to the rest of contents on a global scale. Betweenness is the number of shortest paths between all pairs of nodes in the network that pass through the node under study. This measure also requires global knowledge of the network. It can inform us about which content acts as a bridge between groups of content.
@@ II.1 Official contents of Physics and Mathematics
To illustrate our approach, we focus on the content of two subjects, Physics and Mathematics, from a specific level: the final year of general upper secondary education, “2º de bachillerato”, according to the International Standard Classification of Education Eurostat . In Spain, education is regulated at the national level by an organic law that governs public and private education across primary, secondary, post-secondary and vocational education España (a). Based on this organic law, decrees outlining the minimum educational requirements are established, particularly for secondary education España et al. . These decrees define the basic knowledges, i.e. contents, that each subject must cover. We focus on the contents of Physics and Mathematics. The current curriculum is less centered on content, as the present law emphasizes a competency-based approach Fernández-Gracia et al. (2024), as compared to the previous curriculum that was centered around contents. We rely on the content definitions from the previous curriculum decree España (b), as we created the network before the change in legislation. Nevertheless, both curricula cover the same key concepts with minimal variation in their structure.
The subject of Physics consists of 54 contents distributed in 6 blocks or units, as per the classification of the previous curriculum España (b). Mathematics consists of 36 contents divided into 5 blocks. The titles of the blocks can be read in the Figure 1, which also shows the number of contents for each block.
@@ II.2 Contents network construction
Although this data collection method gathers information on which contents are dependent on others, in the analysis, we treated all relationships as undirected. Specifically, we considered both types 1) and 2) (connection with and without dependency) as equivalent. This decision was made after observing that a majority of the relationships identified during data collection were undirected (type 2), indicating connections between contents that show a certain degree of relatedness but not a strict dependency. Consequently, we replaced the directed relationships (type 1) with undirected ones, disregarding the order in which contents should be taught.
@@ II.3 Identifying central concepts
We investigated 3 different types of centrality. A local measure, which is the degree of each content, i.e. the number of contents related to the same; and two global ones, namely the closeness centrality, which measures how close the rest of the nodes are to one another, and the betweenness centrality, which measures for each content the percentage of all the shortest paths between nodes in the network that pass through one. These centrality measures allow us to create rankings of concepts in terms of importance. The rankings are compared to detect whether there are different central concepts depending on the definition of centrality. The rankings are further compared when the subjects are considered in isolation vs when they are considered together.
@@ II.4 Community structure
The analysis of communities has been done with own code written in python, using the existing implementation of the Infomap algorithm Rosvall et al. (2009); Edler et al. (2020). This algorithm uses random walkers on the network to infer its community structure. A random walker is an agent that can move from one node to another by randomly following the links in the network. If the network has a pronounced community structure, the random walker will remain trapped within each community for a long time before jumping to another. The groupings that arise from the community detection analysis are compared with the grouping into blocks when the subjects are isolated.
@@ III.1 Physics subject
The three centrality measures we calculated (degree, closeness, and betweenness) are highly correlated, meaning that we can rely on just one measure, as the rankings derived from these measures are very similar (the same holds true for Mathematics). This correlation is illustrated in Figure 3. We observe a strictly increasing relationship between degree and closeness centrality, indicating that higher degree values correspond to closer proximity to all other contents. Although the relationship with betweenness centrality is less clear, a notable positive correlation still exists. These correlations collectively support the use of degree as an effective indicator of content importance, without needing the knowledge of the full network structure.
Figure 3: Correlations among the centralities of the three concept networks: A Physics, B Mathematics, and C Physics and Mathematics. The x-axis represents the degree, while the y-axis shows the other two centralities: closeness centrality (blue) and betweenness centrality (orange).
The ten most central contents according to degree can be found in Table 1. The ranking is the same for the centrality of proximity, but it varies minimally for betweenness centrality, with a Spearman rank correlation coefficient of 0.81 ($p<0.001$). The most central content is “Own strategies for scientific activity” in block 1, “Scientific activity”. The next two contents are in block 4 ”Waves” and are related to electromagnetic waves. Among the other 10 central contents, there are three of block 6 “Physics of the XX century”, two of block 4, “Waves”, and two of block 3, “Electromagnetic interaction”. Centralities of all the contents can be checked visually by means of Figure 2.
@@ III.2 Mathematics subject
As in Physics, the correlation between the measures of centrality is high (Figure 3). We see again a perfect match in the rankings by degree and closeness centrality, while that ranking is slightly distorted for betweenness centrality, still with a Spearman rank correlation of 0.88 ($p<0.001$). The 10 most central contents of the network are dominated by contents of the 1st block, “Processes, methods and attitudes in mathematics”, which are very transversal. We also find two contents of the 2nd block, “Numbers and algebra”, related to matrix applications. We can consult these 10 contents in Table 2.
@@ III.3 Physics and Mathematics integration
There is a significant amount of links between the two subjects (red links in Figure 5). It is quite significant that the 5th block of Mathematics, “Statistics and probability”, has more connections with Physics contents than Mathematics. More specifically it is related to quantum physics and its probabilistic interpretation.
As we have seen with the content networks of the subjects separately, all three measures of centrality are also highly correlated by the global network of the two subjects (Figure 3). Again degree and closeness centrality present the same ranking, while the ranking for betweenness centrality correlates with the other with a Spearman rank correlation of 0.95 ($p<0.001$).
Table 3 shows the 10 most central contents of the complete network, ordered according to the degree. The first thing we see is that the list is dominated by transversal contents of the 1st block of the two subjects. We also note that the two contents about electromagnetic waves of the Physics subject continue to be one of the most central. There are two contents that have not appeared before, both from Mathematics subject: “Deductive and inductive reasoning” from Block 1 in the 6th place and “Vectors in three-dimensional space. Scalar, vector and mixed product. Geometrical meaning” from Block 4 in 10th place.
@@
The communities structure has been done again by excluding those contents of the 1st blocks of each subject. The algorithm finds 5 modules. These highly respect the block structure of the contents of the curriculum, grouping some of them. Mathematics Block 2, “Numbers and algebra”, remains as a module by itself. Blocks 2 and 3 of Physics (“Gravitational interaction” and “Electromagnetic interaction”) along with Blocks 3 and 4 of Mathematics (“Analysis” and “Geometry”) form another module, excluding the content “Equations of the line and the plane in space” of Block 4 of Mathematics. This content forms another module together with the contents of Block 5 of Physics, “Geometric optics” and the content “Dispersion. The color” of Block 4 of Physics, “Waves”. The rest of the contents of Block 4 of Physics, “Waves”, forms a module with the contents of Block 6, “Physics of the 20th century”, except for the content “Probabilistic interpretation of quantum physics”. This content forms another module with the contents of Block 5 of Mathematics, “Statistics and probability”.
@@ IV Discussion
Our analysis demonstrates the utility of network theory in organizing and interpreting the relationships within and across educational content. The results show that the three centrality measures—degree, closeness, and betweenness—yielded highly similar results, indicating that the network topology does not require separate centrality criteria for content ordering. This uniformity allows us to rely on degree as a reliable indicator of content importance.
@@ IV.5 Integration of Physics and Mathematics
The integrated Physics and Mathematics network demonstrated a higher average degree compared to the individual networks, driven by numerous interdisciplinary connections. However, the network’s density decreased, reflecting the broader scope of integration. Notable interdisciplinary connections include the dominance of Physics contents like “Electromagnetic Waves” and the emergence of Mathematics contents such as “Vectors in three-dimensional space”. These connections highlight opportunities for integrated STEM education that bridges disciplinary boundaries.
@@ IV.6 Conclusion
This study demonstrates that network analysis provides a robust framework for understanding curricular content. By identifying central and modular contents, this approach supports curriculum designers in prioritizing key topics and exploring interdisciplinary connections. While this work focuses on Physics and Mathematics, the methodology can be extended to other disciplines and educational levels to inform curriculum reform, guide curricular design approach and promote integrated learning experiences by means of teachers cooperation.
@@ VI Data and code availability
All data and code are publicly accessible through the GitHub repository in Fernández-Gracia (2024).

### v3

@@ I Introduction
The application of network science to education is not a novel concept and has been explored in various studies Koponen and Mäntylä (2020). These methods not only facilitate the visualization of complex relationships but also enable the quantification of indirect effects that traditional statistical methods may fail to capture. There are several studies on the coherence of content blocks through the analysis of community structures within students’ concept networks and the centrality of their nodes. Various examples of the use of concept networks are the following: Some studies about general Sayama (2017) or specific topics such as energy and electrostatics Podschuweit and Bernholt (2020); Thurn et al. (2020) use these networks to evaluate and understand students’ ideas and misconceptions. They are also used to evaluate learning processes Siew (2020) or inferential reasoning chains of physics concepts Speirs et al. (2024). In Kubsch et al. (2020), they establish a correlation between the coherence of concept networks and the practical application of knowledge, while in Schwab et al. (2017) teaching sequences are personalized by considering the complex relationships between concepts. Related to concept networks, there are also studies analyzing students’ answers to surveys. For example, in Riihiluoma et al. (2024) the authors analyze students’ understanding of different approaches to quantum mechanics by examining the connections between responses to a questionnaire. Also multiple choice surveys about physics concepts provide interesting information about students’ misconceptions related to other variables Stewart et al. (2021); Wells et al. (2020); Wheatley et al. (2022). Apart from concept networks, network theory in education has been used to evaluate, for example, classmates networks and students’ academic gain Zwolak et al. (2018); Traxler et al. (2018), considering different contexts such as teaching mode (online vs. in-person) Pulgar et al. (2023) or teaching approach (active learning) Traxler et al. (2020). Similarly, authorship networks in publications within the field have also been analyzed Anderson et al. (2017), and social practices between instructors and students have been described using network theory Olsen et al. (2023).
@@ I.1 Basic concepts about complex networks
Centrality measures describe the importance of nodes in a network in different ways. There are many centralities, but in this work we are going to concentrate on three: degree, closeness and betweenness centrality. The degree is the number of links connected to a node and is the most basic centrality measure. It is a local measure that does not require knowledge of the global structure of the network. Typically, a highly connected node is considered very important or central to the network. Nodes with disproportionately large degrees in the network are referred to as hubs. In content networks, nodes with higher degrees represent content that can be related to many other contents, either due to their cross-disciplinary nature, their broad applicability, or because they combine results from different contents. Closeness is the inverse of the average length of the shortest paths that connect a node to the rest of the nodes in the network. In this way, a node that is close to the rest of the nodes will be considered central. This measure requires global knowledge of the network. In content networks, this centrality measure will be able to identify the content that can be more easily related to the rest of contents on a global scale. Betweenness is the number of shortest paths between all pairs of nodes in the network that pass through the node under study. This measure also requires global knowledge of the network. It can inform us about which content acts as a bridge between groups of content. In directed networks, we can measure analogous centralities to those in undirected networks. For degree centrality, we distinguish between in-degree (the number of edges pointing toward a node) and out-degree (the number of edges emerging from a node). Closeness centrality can be assessed in two ways: on the original network, where it measures how easily a node can reach all other nodes, or on the reversed network, where it indicates how easily a node can be reached from any other node. Betweenness centrality naturally extends to directed networks by considering only paths that follow the directionality of edges.
@@ II.1 Official contents of Physics and Mathematics
To illustrate our approach, we focus on the content of two subjects, Physics and Mathematics, as they are displayed in the official curriculum, i.e., the particular law where the basic knowledge or concepts to be taught are established. The relations between these curricular concepts are the subject of our data collection. We focus on a specific level, which is the final year of general upper secondary education, “2º de bachillerato”, according to the International Standard Classification of Education Eurostat . In Spain, education is regulated at the national level by an organic law that governs public and private education across primary, secondary, post-secondary and vocational education España (a). Based on this organic law, decrees outlining the minimum educational requirements are established, particularly for secondary education España et al. . These decrees define the basic knowledges, i.e. contents, that each subject must cover. We focus on the contents of Physics and Mathematics. The current curriculum is less centered on content, as the present law emphasizes a competency-based approach Fernández-Gracia et al. (2024), as compared to the previous curriculum that was centered around contents. We rely on the content definitions from the previous curriculum decree España (b), as we created the network before the change in legislation. Nevertheless, both curricula cover the same key concepts with minimal variation in their structure.
The subject of Physics, according to this curriculum, consists of 54 contents distributed in 6 blocks or units España (b). Mathematics consists of 36 contents divided into 5 blocks. The titles of the blocks can be read in the Figure 1, which also shows the number of contents for each block.
@@ II.2 Contents network construction
We constructed both directed and undirected networks and used one or the other depending on the analysis. In the undirected network, we treated both types 1) and 2) (connections with and without dependency) as equivalent, replacing directed relationships (type 1) with undirected ones and disregarding the order in which contents should be taught. In the directed network, we maintained the original directionality and converted undirected relationships to bidirectional directed edges.
@@ II.3 Identifying central concepts
We investigated three different types of centrality on the undirected networks. A local measure, which is the degree of each content, i.e. the number of contents related to the same; and two global ones, namely the closeness centrality, which measures how close the rest of the nodes are to one another, and the betweenness centrality, which measures for each content the percentage of all the shortest paths between nodes in the network that pass through one. On the directed version of the network we used 5 different variations of the same centralities, namely in-degree, out-degree, closeness on the original and reversed network and betweenness.
These centrality measures allow us to create rankings of concepts in terms of importance. The rankings are compared to detect whether there are different central concepts depending on the definition of centrality. The rankings are further compared when the subjects are considered in isolation vs. when they are considered together.
@@ II.4 Community structure
The analysis of communities has been done on the undirected networks with own code written in Python, using the existing implementation of the Infomap algorithm Rosvall et al. (2009); Edler et al. (2020). This algorithm uses random walkers on the network to infer its community structure. A random walker is an agent that can move from one node to another by randomly following the links in the network. If the network has a pronounced community structure, the random walker will remain trapped within each community for a long time before jumping to another. The groupings that arise from the community detection analysis are compared with the grouping into blocks when the subjects are isolated.
@@ III.1 Physics subject
The eight centrality measures we calculated—degree, in-degree, out-degree, closeness (on the original and reversed networks), and betweenness (for both directed and undirected cases)—are highly correlated, with a minimum Spearman correlation of 0.61 ($p<0.001$, Figure 3). This strong correlation suggests that the rankings derived from these measures are largely consistent, allowing us to rely on degree centrality, the most parsimonious measure, to assess content importance without requiring full knowledge of the network structure.
Directionality does influence the rankings, but not to a great extent. This is confirmed by the fact that, regardless of the specific centrality measure used, at least five concepts are shared between any pair of measures in their respective top ten rankings.
Figure 3: Spearman rank correlation of the pairs of centrality measures on the undirected and directed networks for Physics (A), Mathematics (B) and the joint network (C). The first three centralities are measured on the undirected network: degree (D), closeness (C) and betweenness (B). The other five measures are calculated on the directed version of the networks: in-degree (I-D), out-degree (O-D), closeness on the original and on the reversed directed network (DC and DCR), and betweenness (DB).
The ten most central contents according to degree can be found in Table 1. The most central content is “Own strategies for scientific activity” in block 1, “Scientific activity”. The next two contents are in block 4 “Waves” and are related to electromagnetic waves. Among the other 10 central contents, there are three of block 6 “Physics of the XX century”, two of block 4, “Waves”, and two of block 3, “Electromagnetic interaction”. Centralities of all the contents can be checked visually by means of Figure 2.
@@ III.2 Mathematics subject
The eight centrality measures show a strong and consistent correlation, with the exception of betweenness (both in the directed and undirected cases), which differs slightly from the others (Figure 3). This is reflected in its lower Spearman rank correlation of 0.89 ($p<0.001$) with the remaining measures. Despite this, the overall agreement among centrality rankings remains high, reinforcing the reliability of degree as a representative measure. The consistent relationships between these measures suggest that the network structure inherently supports a shared notion of centrality across different definitions.
Again we will use the degree to rank the nodes, as the most simple measure of centrality. Only betweenness differs in the top ten, but sharing 9 contents. The 10 most central contents of the network are dominated by contents of the 1st block, “Processes, methods and attitudes in mathematics”, which are very transversal. We also find two contents of the 2nd block, “Numbers and algebra”, related to matrix applications. We can consult these 10 contents in Table 2.
@@ III.3 Physics and Mathematics integration
There is a significant amount of links between the two subjects (red links in Figure 5). It is quite significant that the 5th block of Mathematics, “Statistics and probability”, has more connections with Physics contents than Mathematics. More specifically, it is related to quantum physics and its probabilistic interpretation.
As observed in the individual content networks, the eight centrality measures in the global network of both subjects remain highly correlated (Figure 3). The results reflect a mixture of those obtained for the separate networks, with some variation in the correlations between centrality measures. However, these variations are less pronounced than in the Physics network, with a minimum Spearman rank correlation of 0.91 ($p<0.001$). This consistency further supports the robustness of the centrality rankings across different network structures.
Table 3 shows the 10 most central contents of the complete network, ordered according to the degree. When using other centrality measures the top ten nodes remain the same except for betweenness in the undirected case. Still, this ranking retains 9 of the top ten nodes. The first thing we see is that the list is dominated by transversal contents of the 1st block of the two subjects. We also note that the two contents about electromagnetic waves of the Physics subject continue to be one of the most central. There are two contents that have not appeared before, both from the Mathematics subject: “Deductive and inductive reasoning” from Block 1 in the 6th place and “Vectors in three-dimensional space. Scalar, vector and mixed product. Geometrical meaning” from Block 4 in 10th place.
@@
The communities structure has been done again by excluding those contents of the 1st blocks of each subject. The algorithm finds 5 modules. These highly respect the block structure of the contents of the curriculum, grouping some of them. Mathematics Block 2, “Numbers and algebra”, remains as a module by itself. Blocks 2 and 3 of Physics (“Gravitational interaction” and “Electromagnetic interaction”) along with Blocks 3 and 4 of Mathematics (“Analysis” and “Geometry”) form another module, excluding the content “Equations of the line and the plane in space” of Block 4 of Mathematics. This content forms another module together with the contents of Block 5 of Physics, “Geometric optics” and the content “Dispersion. The color” of Block 4 of Physics, “Waves”. The rest of the contents of Block 4 of Physics, “Waves”, form a module with the contents of Block 6, “Physics of the 20th century”, except for the content “Probabilistic interpretation of quantum physics”. This content forms another module with the contents of Block 5 of Mathematics, “Statistics and probability”.
@@ IV Discussion
Our analysis demonstrates the utility of network theory in organizing and interpreting the relationships within and across educational content. The results show that the eight centrality measures yielded highly similar results, indicating that the network topology does not require separate centrality criteria for content ordering. This uniformity allows us to rely on degree as a reliable indicator of content importance, regardless also of directionality.
@@ IV.1 Centrality Insights
The primary purpose of Physics in the final year of high school is to connect the limitations of classical models with the basic aspects (at least qualitatively) of modern physics. This is consistent with the distribution of the most central concepts shown in the Table 1. In addition to the procedural aspect (“Specific strategies to scientific activity”), these are essentially: 1) waves, 2) fields, and 3) issues related to the introduction of modern physics. Concerning point 3, we find: fundamental interactions (beyond gravity and electromagnetism, i.e., the strong and weak interactions), quantum physics, and a topic called “limitations of classical models,” which leads to the new insights of 20th-century physics. Waves, the concept of fields, and the general aspects of modern physics serve as the core around which the other concepts revolve in a pre-university physics course.
In the case of Mathematics, the most central concepts are mostly procedural (from the first block). The same happens in the joint network, where the most central concepts are dominated by the procedural contents of Mathematics, as we will discuss below. And those of Physics that either relate to the procedural part or deal with waves.
@@ IV.5 Integration of Physics and Mathematics
The integrated Physics and Mathematics network demonstrated a higher average degree compared to the individual networks, driven by numerous interdisciplinary connections. However, the network’s density decreased, reflecting the broader scope of integration. Notable interdisciplinary connections include the dominance of Physics contents like “Electromagnetic Waves” and the emergence of Mathematics contents such as “Vectors in three-dimensional space”. These connections highlight opportunities for integrated STEM education that bridge disciplinary boundaries.
@@ IV.6 Conclusion
This study demonstrates that network analysis provides a robust framework for understanding curricular content. By identifying central and modular contents, this approach supports curriculum designers in prioritizing key topics and exploring interdisciplinary connections. While this work focuses on Physics and Mathematics, the methodology can be extended to other disciplines and educational levels to inform curriculum reform, guide curricular design approaches, and promote integrated learning experiences by means of teachers’ cooperation.
@@ VI Data and code availability
All data and code are publicly accessible through the GitLab repository in Fernández-Gracia (2024)